OpenStudy (anonymous):

Find the product. (-7p)^3 i dont get this one..

OpenStudy (anonymous):

\[(-7)^3 \times p^3\] Solve this..

OpenStudy (anonymous):

-343p^3?

OpenStudy (anonymous):

\[(-7) \times (-7) \times (-7) = \]

OpenStudy (anonymous):

Yes absolutely right..

OpenStudy (usukidoll):

ok so we have an odd exponent and a negative number. Whenever we have this issue the ending result will always be negative (-7)^3 means -7 x -7 x -7 = it will become -343 -7 x -7 = 49 x - 7 = -343 hence the negative sign.. Remember that a negative x negative = positive. the (p)^3 is simple. the exponent determines the number of times a letter (variable) must be written. It's like writing p 3 times p p p now combine them -343p^3

OpenStudy (anonymous):

\[(-7p)^3 = (-7)^3+(p)^3\]\[(-7)\times(-7)\times(-7)= -343\]\[(p)\times(p)\times(p)= p^3\]\[-343+p^3 = -343p^3\] \[(-7p)^3=-343p^3\]

OpenStudy (anonymous):

^^ i hope this helps! :)